# number of leading zeros ICWS improved consistent weighted sampling

### v1

@@ 2.5. Saving Random Values
Algorithm 4 requires many random values. For every interval $\mathcal{I}_{m}$ a bit $X_{m}\in\{0,1\}$, an integer $Y_{m}\in\mathcal{I}_{m}$, and an integer sequence $(Z_{m,v})_{v\geq 0}$ with $Z_{m,v}\in\llbracket 2^{m+1}\rrbracket$ need to be sampled. We are free to choose in which order those random values are generated. However, the order must be well-defined and not depend on $n$ in order to preserve monotonicity.
@@
$m_{i+1}=\lfloor\log_{2}(U\bmod 2^{m_{i}})\rfloor$
@@
$m_{0}=\lfloor\log_{2}(n-1)\rfloor+1$
as $\mathcal{I}_{m_{0}}$ is the smallest interval without overlap with $\llbracket n\rrbracket$.
Among the set of intervals $\{\mathcal{I}_{m}\}_{m\in\llbracket M\rrbracket}$, at most two need to be processed as $\mathcal{I}_{{m_{1}}}$ and $\mathcal{I}_{{m_{2}}}$ are the largest two intervals that contain at least one active index and overlap with $\llbracket n\rrbracket$. As $\mathcal{I}_{{m_{2}}}\subseteq\llbracket n\rrbracket$ by definition, $\mathcal{I}_{{m_{2}}}$ must contain an active index that is also smaller than $n$. Therefore, the active indices of smaller intervals $\mathcal{I}_{m}$ with $m<{m_{2}}$ are not relevant. Furthermore, within $\mathcal{I}_{{m_{2}}}$ just the largest active index is needed, which means that the random sequence $(Z_{{m_{2}},v})_{v\geq 0}$ is never used. The only sequence that may be required for generating the largest active index smaller than $n$ is $(Z_{{m_{1}},v})_{v\geq 0}$ as the active indices of $\mathcal{I}_{m_{1}}$ are not necessarily smaller than $n$.
@@
$Z_{m,v}=W_{v}\bmod 2^{m+1}.$
It should be noted that the generation of active indices is no longer statistically equivalent to (4), even if this modification preserves uniformity and monotonicity.
@@
$Y_{m}=2^{m}+(V_{s}\bmod 2^{m})\qquad\text{with $\textstyle s=(\sum_{i=0}^{m}X_{i})\bmod 2$}$
@@
In summary, it is sufficient to generate $V_{0}$ and $V_{1}$ followed by the sequence $(W_{v})_{v\geq 0}$ to generate all needed random values without breaking uniformity. Therefore, all these values can be taken from a single sequence which allows to use a standard PRG.
@@ 2.6. JumpBackHash
       while $u\neq 0$ do
@@
              $\triangleright$ equals $m_{i}$ in $i$-th iteration
@@
             for $v\leftarrow 0$ to $\infty$ do
@@
                    $\triangleright$ satisfied latest in $2^{\text{nd}}$ while-iteration $W_{v}\leftarrow\mathcal{R}[\operatorname*{Uniform}(\llbracket 2^{M}\rrbracket)]$
@@
              $\triangleright$ has same effect as $u\leftarrow u\bmod 2^{m}$
@@
Applying all these optimizations to Algorithm 4 finally leads to JumpBackHash as shown in Algorithm 5. The variables $U$ and $W_{v}$ as well as the loop variable $v$ could be eliminated, but are kept for better comprehensibility and better comparison with Algorithm 4.
Algorithm 6 shows how the algorithm can be implemented with a few lines of code in Java. Expressions of kind $\lfloor\log_{2}u\rfloor$ can be replaced by $31-\operatorname*{nlz}(u)$ if $u$ is a 32-bit integer. Furthermore, bit-shift operations are well-defined in Java. For example, shifting by $-29$ bits is the same as shifting by 3 bits. Therefore, shifting by $31-\operatorname*{nlz}(u)$ is the same as shifting by $-1-\operatorname*{nlz}(u)={\sim}\operatorname*{nlz}(u)$ bits where $\sim$ denotes the bitwise complement. The number of leading zeros (NLZ) and also the bitcount operations can usually be translated directly into specific CPU instructions and are therefore very fast.
@@ 2.7. Runtime Analysis
First, we need two random values for $V_{0}$ and $V_{1}$ (compare Section 2.5). Further random values are only needed if the largest active index in $\mathcal{I}_{m_{1}}$ is greater than or equal to $n$. This can only happen, if $m_{1}=m_{0}-1$. The probability that $\mathcal{I}_{m_{0}-1}$ contains an active index at all is $\frac{1}{2}$. As the largest active index would be uniformly distributed over $\mathcal{I}_{m_{0}-1}$, the probability that $\mathcal{I}_{m_{0}-1}$ contains an active index that is larger than or equal to $n$ is given by $\frac{1}{2}\cdot\frac{2^{m_{0}}-n}{2^{m_{0}-1}}=1-\alpha_{n}^{-1}$ with
@@
Only if this is the case we need to start generating the random sequence $(W_{v})_{v\geq 0}$ to find the largest active index within $\mathcal{I}_{m_{0}-1}$ that is actually smaller than $n$. The inner loop is left as soon as $W_{v}\bmod 2^{m_{1}+1}=W_{v}\bmod 2^{m_{0}}<n$ which occurs with probability $\frac{n}{2^{m_{0}}}=\alpha_{n}^{-1}$. Therefore, the number of random values consumed within the inner loop follows a geometric distribution with positive support and a success probability of $\alpha_{n}^{-1}$.
@@
(8) $T_{\text{JBH}}\sim 2+\operatorname*{Bernoulli}(1-\alpha_{n}^{-1})\cdot\operatorname*{Geom}(\alpha_{n}^{-1}).$
Using the identities $\operatorname*{\mathbb{E}}(XY)=\operatorname*{\mathbb{E}}(X)\operatorname*{\mathbb{E}}(Y)$ and $\operatorname*{Var}(XY)=(\operatorname*{\mathbb{E}}(X))^{2}\operatorname*{Var}(Y)+(\operatorname*{\mathbb{E}}(Y))^{2}\operatorname*{Var}(X)+\operatorname*{Var}(X)\operatorname*{Var}(Y)$ for independent $X$ and $Y$ together with $\operatorname*{\mathbb{E}}(\operatorname*{Bernoulli}(p))=p$, $\operatorname*{Var}(\operatorname*{Bernoulli}(p))=p(1-p)$, $\operatorname*{\mathbb{E}}(\operatorname*{Geom}(p))=\frac{1}{p}$, and $\operatorname*{Var}(\operatorname*{Geom}(p))=\frac{1-p}{p^{2}}$, the expectation and the variance of $T_{\text{JBH}}$ can be derived as
@@
(11) $T_{\text{JBH*}}\sim 1+\operatorname*{Bernoulli}(1-\alpha_{n}^{-1})\cdot\operatorname*{Geom}(\alpha_{n}^{-1}(2-\alpha_{n}^{-1}))$
with expectation and variance given by
@@ 3. Experimental Verification
To validate JumpBackHash, we have performed a series of tests. We used the SplitMix random generator (Steele et al., [n.d.]) which produces a sequence of 64-bit random integers, is known to have good statistical properties (Lemire, [n.d.]), and is also available in the Java standard library as SplittableRandom. The source code to run all tests and to reproduce the presented results has been published at https://github.com/dynatrace-research/jumpbackhash-paper. The tested JumpBackHash implementation is equivalent to the one included in our Hash4j Java open source library since version 0.17.0 (Has, [n.d.]).
Figure 1. The measured mean and variance of the number of consumed random values for JumpHash and JumpBackHash. JumpBackHash* corresponds to the case when a single random value is used for two as in Algorithm 6.
@@ 3.1. Consistency
To verify consistency we must verify monotonicity and uniformity. Monotonicity is straightforward to test. We did 10k simulation runs, in each of which we used a random key and computed the assignments for all $n\in\{1,2,\ldots,10000\}$ and checked, whenever the reassignment changed after an increment of $n$, that the new assignment corresponds to the new bucket according to (3).
Testing uniformity is more sophisticated as it requires statistical means. Our strategy was to sample many bucket indices by calling the consistent hash function for many different keys. In our test we always used 1 million randomly generated keys. If uniformity holds, we expect that the bucket index distribution is uniform. For small numbers of buckets, as long as it is much smaller than the number of keys, this can be verified by applying the G-test (McDonald, 2014). In this way we checked the uniformity for all $n\in\{1,2,\ldots,1000\}$. Additionally, we also investigated if uniformity holds for large $n$ close to $12^{31}-$ 1 which is the largest possible positive 32-bit integer value in Java. In particular, we have tested the values $n\in\{2^{31}-1,2^{31}-2,2^{30}+1,2^{30},2^{30}-1,3\cdot 2^{28},2^{29}+1,2^{29},2^{29}-1,3\cdot 2^{27},2^{28}+1,2^{28},2^{28}-1\}$ using the Kolmogorov-Smirnov test (Kanji, 2006), as the number of occurrences per possible outcome would be too low for a valid application of the G-test.
@@ 3.2. Random Value Consumption
This results in 7482 integer values that are roughly uniformly distributed on the logarithmic scale over the range $[1,10^{6}]$.
@@ 3.3. Performance Measurements
The number of buckets was varied from 1 to 1 million. We included all values that can be represented as $2^{i},2^{i}+1,\lfloor 2^{i}(1+\frac{1}{4})\rfloor,\lfloor 2^{i}(1+\frac{1}{2})\rfloor,\lfloor 2^{i}(1+\frac{3}{4})\rfloor$ with some integer $i$. The first two correspond to the best and the worst cases for JumpBackHash according to our theoretical analysis.
The benchmark results for JumpHash (Algorithm 2) and JumpBackHash* (Algorithm 6) are shown in Figure 2. As reference, we also included non-consistent approaches based on the modulo-operation (1) and a pseudorandom assignment using Lemire’s algorithm (Lemire, 2019). Furthermore, we added a dummy test that just consumed the random keys to make sure that the input is not the bottleneck. To allow a fairer comparison, we used SplitMix (Steele et al., [n.d.]) for generating random values for all approaches that require a PRG.
@@
Figure 2. Benchmark results on an Amazon EC2 c5.metal instance with Intel Xeon Platinum 8275CL CPU and processor P-state set to 1.
@@ 4. Future Work
In our tests we used the SplitMix random generator (Steele et al., [n.d.]) which is already quite fast and known to have good quality (Lemire, [n.d.]). However, there might be even faster random generators such as linear congruential random generators (L’Ecuyer, [n.d.]) that have sufficient statistical quality for JumpBackHash.

### v2

@@ 2.5. Saving Random Values
Algorithm 4 requires many random values. For every interval $\mathcal{I}_{m}$ a bit $X_{m}\in\{0,1\}$, an integer $Y_{m}\in\mathcal{I}_{m}$, and an integer sequence $(Z_{m,v})_{v\geq 0}$ with $Z_{m,v}\in\llbracket 2^{m+1}\rrbracket$ need to be sampled when using option 2. The sequences $(Z_{m,v})_{v\geq 0}$ are finite, as it is sufficient to generate values up to the first value that is smaller than $2^{m}$ and that causes leaving the inner loop of Algorithm 4 in any case. We are free to change the order in which those independent random values and sequences are generated using the same PRG without changing the statistical properties of Algorithm 4. However, the order must be well-defined and not depend on $n$ in order to preserve monotonicity.
@@
(15) $m_{i+1}=\lfloor\log_{2}(U\bmod 2^{m_{i}})\rfloor$
@@
(16) $m_{0}=\lfloor\log_{2}(n-1)\rfloor+1\qquad\text{for $n\geq 2$}$
which is the solution of the inequality
$2^{m_{0}-1}<n\leq 2^{m_{0}}.$
Thus, $\mathcal{I}_{m_{0}}$ is the smallest interval without overlap with $\llbracket n\rrbracket$.
Among the set of intervals $\{\mathcal{I}_{m}\}_{m\in\llbracket M\rrbracket}$, at most two need to be processed as $\mathcal{I}_{{m_{1}}}$ and $\mathcal{I}_{{m_{2}}}$ (if they exist) are the largest two intervals that contain at least one active index and overlap with $\llbracket n\rrbracket$. As $\mathcal{I}_{{m_{2}}}\subseteq\llbracket n\rrbracket$ by definition, $\mathcal{I}_{{m_{2}}}$ must contain an active index that is also smaller than $n$. Therefore, the active indices of smaller intervals $\mathcal{I}_{m}$ with $m<{m_{2}}$ are not relevant. Furthermore, within $\mathcal{I}_{{m_{2}}}$ just the largest active index is needed, which means that the random sequence $(Z_{{m_{2}},v})_{v\geq 0}$ is never used. The only sequence that may be required for generating the largest active index smaller than $n$ is $(Z_{{m_{1}},v})_{v\geq 0}$ as the active indices of $\mathcal{I}_{m_{1}}$ are not necessarily smaller than $n$.
@@
(17) $Z_{m,v}=W_{v}\bmod 2^{m+1}.$
It should be noted that with this modification, the generation of active indices is no longer statistically equivalent to (4). However, uniformity is preserved as the consumed random values are still independent. Monotonicity is also still ensured, as the active indices are still generated in descending order.
@@
(18) $Y_{m}=2^{m}+(V_{s}\bmod 2^{m})\qquad\text{with $\textstyle s=(\sum_{i=0}^{m}X_{i})\bmod 2$}$
@@
In summary, it is sufficient to generate $V_{0}$ and $V_{1}$ followed by the sequence $(W_{v})_{v\geq 0}$ to generate all needed random values without breaking uniformity and monotonicity. Therefore, all these values can be taken from a single random sequence which allows to use a standard PRG.
@@ 2.6. JumpBackHash
       while $u\neq 0$ do $\triangleright$ iterate over intervals $\mathcal{I}_{m_{1}},\mathcal{I}_{m_{2}},\ldots$ that contain at least one active index
@@
              $\triangleright$ equals $m_{i}$ in $i$-th iteration, compare (15)
@@
              $\triangleright$ corresponds to $Y_{m}$, compare (18)
             for $v\leftarrow 0$ to $\infty$ do $\triangleright$ $v$ could be eliminated but is kept for better comparability with Algorithm 4
@@
                    $\triangleright$ condition is always satisfied latest in the second while-iteration $W_{v}\leftarrow\mathcal{R}[\operatorname*{Uniform}(\llbracket 2^{M}\rrbracket)]$
@@
                    $\triangleright$ stop if there are no further active indices in $\mathcal{I}_{m}$
@@
              $\triangleright$ has the same effect as $u\leftarrow u\bmod 2^{m}$
@@
Applying all these optimizations to Algorithm 4 with option 2 finally leads to JumpBackHash as shown in Algorithm 5. The variables $U$ and $W_{v}$ as well as the loop variable $v$ could be eliminated, but are kept for better comprehensibility and better comparability with Algorithm 4. The while loop iterates over the intervals $\mathcal{I}_{m_{1}},\mathcal{I}_{m_{2}},\ldots$ according to (15). Since we know that $\mathcal{I}_{m_{2}}$, if it exists, must contain an active index smaller than $n$, the condition for the return statement within the inner loop for $\mathcal{I}_{m_{2}}$ is always fulfilled. This means that there will never be more than two iterations of the while loop.
Algorithm 6 shows how the algorithm can be implemented with a few lines of code in Java. Expressions of kind $\lfloor\log_{2}u\rfloor$ can be replaced by 31-numberOfLeadingZeros(u) if $u$ is a positive 32-bit integer. Furthermore, bit-shift operations are well-defined in Java. For example, shifting by $-29$ bits is the same as shifting by 3 bits. Therefore, shifting by 31-numberOfLeadingZeros(u) is the same as shifting by -1-numberOfLeadingZeros(u) or by ~numberOfLeadingZeros(u) when using Java’s bitwise complement operator. The numberOfLeadingZeros and bitCount Java functions are intrinsic candidates. They are usually translated directly into specific CPU instructions and are therefore very fast.
@@ 2.7. Runtime Analysis
First, we need two random values for $V_{0}$ and $V_{1}$ (compare Section 2.5). Further random values are only needed if the largest active index in $\mathcal{I}_{m_{1}}$ is greater than or equal to $n$. This can only happen, if $n\in\mathcal{I}_{m_{1}}$, and thus if $m_{1}=m_{0}-1$. The probability that $\mathcal{I}_{m_{0}-1}$ contains an active index at all is $\frac{1}{2}$. As the largest active index would be uniformly distributed over $\mathcal{I}_{m_{0}-1}$, the probability that $\mathcal{I}_{m_{0}-1}$ contains an active index that is larger than or equal to $n$, or equivalently from $[n,2^{m_{0}})$, is given by $\frac{1}{2}\cdot\frac{2^{m_{0}}-n}{2^{m_{0}-1}}=1-\alpha_{n}^{-1}$ with
@@
Only if this is the case we need to start generating the random sequence $(W_{v})_{v\geq 0}$ to find the largest active index within $\mathcal{I}_{m_{0}-1}$ that is actually smaller than $n$. The inner loop is left, either via the break or the return statement, as soon as $W_{v}\bmod 2^{m_{1}+1}=W_{v}\bmod 2^{m_{0}}<n$ which occurs with probability $\frac{n}{2^{m_{0}}}=\alpha_{n}^{-1}$. Therefore, the number of random values consumed within the inner loop follows a geometric distribution with positive support and a success probability of $\alpha_{n}^{-1}$.
@@
(21) $T_{\text{JBH}}\sim 2+\operatorname*{Bernoulli}(1-\alpha_{n}^{-1})\cdot\operatorname*{Geom}(\alpha_{n}^{-1})\qquad\text{for $n\geq 2$}.$
Using the identities for the mean and variance of the product of independent $K$ and $L$ (Goodman, 1960)
$\operatorname*{\mathbb{E}}(KL)$ $=\operatorname*{\mathbb{E}}(K)\operatorname*{\mathbb{E}}(L),$
$\operatorname*{Var}(KL)$ $=(\operatorname*{\mathbb{E}}(K))^{2}\operatorname*{Var}(L)+(\operatorname*{\mathbb{E}}(L))^{2}\operatorname*{Var}(K)+\operatorname*{Var}(K)\operatorname*{Var}(L)$
together with (8), (9), and the known expressions for geometrically distributed random variables
$\operatorname*{\mathbb{E}}(\operatorname*{Geom}(p))$ $=\frac{1}{p},$
$\operatorname*{Var}(\operatorname*{Geom}(p))$ $=\frac{1-p}{p^{2}},$
the mean and the variance of $T_{\text{JBH}}$ can be derived as
@@
(24) $T_{\text{JBH*}}\sim 1+\operatorname*{Bernoulli}(1-\alpha_{n}^{-1})\cdot\operatorname*{Geom}(\alpha_{n}^{-1}(2-\alpha_{n}^{-1}))\qquad\text{for $n\geq 2$}$
with mean and variance given by
@@
$\alpha_{n}$ is minimized and equal to 1 if $n$ is exactly a power of two. The maxima of $\alpha_{n}$ occur whenever the number of buckets are powers of two incremented by 1, $n=2^{i}+1$ and approach 2 as $n\rightarrow\infty$. This leads to a sawtooth pattern on the log-scale. As the expressions (22), (23), (25), and (26) are all monotonically increasing functions of $\alpha_{n}$ over $[1,2)$, they replicate this sawtooth pattern. In contrast to JumpHash, where not only the mean (10) but also the variance (11) increases logarithmically, both are bounded in case of JumpBackHash.
@@ 3. Experimental Verification
To validate JumpBackHash, we have performed a series of tests. We used the SplitMix random generator (Steele et al., 2014) which produces a sequence of 64-bit random integers, is known to have good statistical properties (Lemire, [n. d.]), and is also available in the Java standard library as SplittableRandom. The source code to run all tests and to reproduce the presented results has been published at https://github.com/dynatrace-research/jumpbackhash-paper. The tested JumpBackHash implementation is equivalent to the one included in our Hash4j Java open source library since version 0.17.0 (Has, [n. d.]).
Figure 2. The measured mean and variance of the number of consumed random values for JumpHash and JumpBackHash. JumpBackHash* corresponds to the case when a single random value is split into two as in Algorithm 6.
@@ 3.1. Consistency
[…]
Testing uniformity is more sophisticated as it requires statistical means. Our strategy was to sample many bucket indices by calling the consistent hash function for many different keys. In our test we always used 1 million randomly generated keys. If uniformity holds, we expect that the bucket index distribution is uniform. For small numbers of buckets, as long as it is much smaller than the number of sampled keys, this can be verified by applying the G-test (McDonald, 2014). In this way we checked the uniformity for all $n\in\{1,2,\ldots,$1000$\}$. Additionally, we also investigated if uniformity holds for large $n$ close to $12^{31}-$ 1 which is the largest possible positive 32-bit integer value in Java. In particular, we have tested the values $n\in\{2^{31}-1,2^{31}-2,2^{30}+1,2^{30},2^{30}-1,3\cdot 2^{28},2^{29}+1,2^{29},2^{29}-1,3\cdot 2^{27},2^{28}+1,2^{28},2^{28}-1\}$ using the Kolmogorov-Smirnov test (Kanji, 2006), as the number of occurrences per possible outcome would be too low for a valid application of the G-test.
@@ 3.2. Random Value Consumption
This results in 7482 positive integer values that are roughly uniformly distributed on the logarithmic scale over the range $[1,10^{6}]$.
@@ 3.3. Performance Measurements
Figure 3. Benchmark results on an Amazon EC2 c5.metal instance with Intel Xeon Platinum 8275CL CPU and processor P-state set to 1. Unlike Figure 2, a logarithmic scale is used for the y-axis.
@@
[…]
The benchmark results for JumpHash (Algorithm 2) and JumpBackHash* (Algorithm 6) are shown in Figure 3. As reference, we also included non-consistent approaches based on the modulo-operation (1) and a pseudorandom assignment using Lemire’s algorithm (Lemire, 2019). Furthermore, we added a dummy test that just consumed the random keys to make sure that the input is not the bottleneck. To allow a fairer comparison, we used SplitMix (Steele et al., 2014) for generating random values for all approaches that require a PRG.
@@ 4. Future Work
In our tests we used the SplitMix random generator (Steele et al., 2014) which is already quite fast and known to have good quality (Lemire, [n. d.]). However, there might be even faster random generators such as linear congruential random generators (L’Ecuyer, 1999) that have sufficient statistical quality for JumpBackHash.